OpenStudy (peterg2001):

help please will medal The following table shows a direct variation. 3 15 18 27 y 5 6 9 Find y. 0 1 2 3 @danjs

OpenStudy (peterg2001):

@bea

OpenStudy (peterg2001):

@hhelpplzzzz

OpenStudy (anonymous):

1

OpenStudy (lizz123):

You divide by 3 each one 15/3= 5 18/3=6 27/3=9 3/3=1 y=1

OpenStudy (danjs):

15 is to 5 as 3 is to y

OpenStudy (danjs):

or you can do the y=k*x thing, and use one point (x,y) to figure out k